the equivalent weight of a metal is 36 . what weight of the metal would give 9.322 gm of its chloride

Grade:12

2 Answers

Akash Kumar Dutta
10 years ago

x/36=9.322/36+35.5
find x

BAYANA SAGAR
10 years ago

x/36=9.322/36+35.

=1269.322